What is the mass in grams of potassium ion K+

(30.0 g/mol) in 25.0 g of potassium 

sulfate (K2SO4)?


1
Expert's answer
2021-03-25T04:10:16-0400

Solution:

The molecular weight of potassium ion (K+) is 39.10 g/mol (not 30.0 g/mol)

K2SO4 → 2K+ + SO42-

(25.0 g K2SO4) × (1 mol K2SO4 / 174.26 g K2SO4) × (2 mol K+ / 1 mol K2SO4) × (39.10 g K+ / 1 mol K+) = 11.2188 g K+ = 11.2 g K+

Mass of K+ = 11.2 g


OR (another solution path):

%K+ = (2 × Molar mass of K+ × 100%) / Molar mass of K2SO4

%K+ = (2 × 39.10 g mol-1 × 100%) / 174.26 g mol-1 = 44.88%

Mass of K+ = (%K+ × Mass of K2SO4) / 100%

Mass of K+ = (44.88% × 25.0 g) / 100% = 11.22 g = 11.2 g

Mass of K+ = 11.2 g


Answer: The mass of potassium ion (K+) is 11.2 g.